OpenStudy (anonymous):

(x - 5)2 + (y + 9)2 = 144 Which statement is true? The circle is centered at (-5, -9) and has a diameter of 12. The circle is centered at (5, -9) and has a radius of 12. The circle is centered at (5, 9) and has a diameter of 12. The circle is centered at (5, 9) and has a radius of 12.

OpenStudy (aravindg):

The equation of a circle is \[(x-h)^2+(y-k)^2=r^2\] where (h,k) is centre and r is radius .Can you answer it now?

OpenStudy (anonymous):

So would it be C?

OpenStudy (aravindg):

nope?Try again

OpenStudy (jim766):

it will be the opposite of each number in the parenthese

OpenStudy (anonymous):

D? The original is (x - 5)2 + (y + 9)2 = 144 so I thought maybe 12 was the diameter

OpenStudy (aravindg):

nooo!

OpenStudy (anonymous):

Okay then I'm confused

OpenStudy (aravindg):

Its r^2 on right 144=12^2 12 is radius!!

OpenStudy (jim766):

opposite of -5 is 5 opposite of 9 is -9

OpenStudy (aravindg):

Its r^2!!

OpenStudy (anonymous):

So it is D

OpenStudy (aravindg):

nope

OpenStudy (jim766):

(5, -9)

OpenStudy (aravindg):

notice there is a difference here from general equation

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Y'all both are confusing me lol

OpenStudy (aravindg):

Its (x+9)=(x-(-9) Hence center will be (5,-9) and radius 12

OpenStudy (aravindg):

That will be B

OpenStudy (anonymous):

Okay can you help me with one more?

OpenStudy (aravindg):

Yes. Post it separately.